OpenStudy (-welp-):

How is sqrt20 not in simplified radical form?

zepdrix (zepdrix):

Because 20 contains a factor which is a `perfect square`! :)

zepdrix (zepdrix):

If you break down 20 into its prime factorization, it might become more clear,\[\large\rm 20=4\cdot5=2\cdot2\cdot5\]

zepdrix (zepdrix):

\[\large\rm \sqrt{20}=\sqrt{4\cdot5}=\sqrt{4}\cdot\sqrt{5}\]Square root of 4 can be simplified, ya?